### v1

# Admissibility Breakdown in High-Dimensional Sparse Regression with $\ell_{1}$ Regularization

Abstract
The choice of the tuning parameter in the Lasso is central to its statistical performance in high-dimensional linear regression. Classical consistency theory identifies the rate of the Lasso tuning parameter, and numerous studies have established non-asymptotic guarantees. Nevertheless, the question of optimal tuning within a non-asymptotic framework has not yet been fully resolved. We establish tuning criteria above which the Lasso becomes inadmissible under mean squared prediction error. More specifically, we establish thresholds showing that certain classical tuning choices yield Lasso estimators strictly dominated by a simple Lasso–Ridge refinement. We also address how the structure of the design matrix and the noise vector influences the inadmissibility phenomenon.
1 Introduction
The Lasso (Tibshirani, 1996) is one of the most widely used methods for high-dimensional linear regression. By combining estimation and variable selection through an $\ell_{1}$ penalty, it leads to computationally efficient procedures with solid theoretical properties. A large body of work has established oracle inequalities, prediction error bounds, and sparsity recovery guarantees under various assumptions on the design matrix; see, for example, (van de Geer, 2016) and the references therein. In this literature, it is standard to choose the tuning parameter at the rate $\lambda\asymp\sigma\sqrt{\log(p)/n}$, which ensures favorable prediction behavior in high-dimensional settings.
More recently, several studies have suggested that the choice of the tuning parameter can be more delicate, especially when the design exhibits strong correlations. For instance, (van de Geer and Lederer, 2013; Hebiri and Lederer, 2013; Dalalyan et al., 2017) show that tuning parameters substantially smaller than the classical choice may yield sharp oracle inequalities and, in some cases, improved prediction performance. Related results in (Hebiri and Lederer, 2013) indicate that stronger correlations in the design matrix call for smaller tuning parameters to obtain sharper bounds. In a different direction, (Bellec, 2018) introduced the notion of a “noise barrier,” showing that the minimal penalty level may depend on intrinsic quantities such as sparsity, leading to refined scales of order $\lambda\asymp\sigma\sqrt{\log(p/k)/n}$ rather than $\sigma\sqrt{\log(p)/n}$. Taken together, these works provide theoretical support for the empirical observation that tuning parameters smaller than the classical rate can be advantageous in certain regimes. However, these results are primarily expressed in terms of refined risk bounds and do not directly address whether the Lasso equipped with the traditional tuning parameter can be improved upon in a uniform sense.
In this paper, we provide a decision-theoretic perspective on Lasso tuning and prediction improvement. Rather than focusing on refined prediction bounds with new tuning choices, we ask whether the Lasso, for a given tuning parameter, can be dominated in mean squared prediction error by a suitably constructed refinement, as proposed in a previous arXiv preprint (Liu, 2025). We show that, over a broad range of tuning values, including the standard Lasso tuning, the Lasso estimator is dominated in prediction risk by an Lasso–Ridge refinement. Although the Lasso serves as our primary example, the analysis extends, to some extent, to other sparse estimators induced by $\ell_{1}$ regularization. Our results provide a complementary explanation for the empirical finding that tuning parameters smaller than the standard theoretical choice often yield superior prediction performance. This perspective also highlights a distinction between rate-based guarantees and uniform risk improvement, and clarifies how tuning choices interact with sparsity and design structure.
The main contributions of this paper can be summarized as follows:
  1. 1.
    We introduce a framework for comparing prediction risks that allows direct assessment of whether a given Lasso tuning choice can be improved upon.
  2. 2.
    We establish criteria for Lasso tuning parameters beyond which the Lasso is dominated in mean squared prediction error, showing that a range containing commonly used tuning parameters can lead to suboptimal performance.
  3. 3.
    We provide insight into how intrinsic structural features, such as sparsity and properties of the design matrix, affect the inadmissibility phenomenon.
2 Preliminaries and Background
2.1 Notation and Model Assumptions
The cardinality of a set $E$ is denoted by $|E|$. For a vector $x$, define $\|x\|_{1}=\sum_{i}|x_{i}|$ and $\|x\|_{2}=\bigl(\sum_{i}x_{i}^{2}\bigr)^{1/2}$. For a matrix $A=(A_{ij})\in\mathbb{R}^{n\times p}$, define
$\|A\|_{\infty}$ $=\max_{i}\sum_{j}|A_{ij}|,$
$\|A\|_{2}$ $=\sup_{\|x\|_{2}=1}\,\|Ax\|_{2}.$
For any $S\subseteq[p]$, $A_{S}$ denotes the submatrix consisting of the columns of $A$ indexed by $S$, and $\beta_{S}$ denotes the corresponding subvector of $\beta$. We adopt the standard conventions for zero-dimensional vectors and matrices. In particular, a $0\times n$ matrix and a $0\times 1$ vector are well defined, and the $0\times 0$ matrix is treated as a valid square matrix whose inverse is itself. Under these conventions, matrix multiplication involving zero-dimensional objects is well defined and dimensionally consistent. We work with the linear model
$y=X\beta^{0}+\varepsilon,$
where $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times p}$, and $\beta^{0}\in\mathbb{R}^{p}$. The noise vector $\varepsilon\in\mathbb{R}^{n}$ is arbitrary and no specific distributional assumption is imposed unless stated otherwise. Columns of $X$ are normalized so that $\|X_{j}\|_{2}^{2}=n$ for all $j=1,\dots,p$.
2.2 Lasso and Ridge: Definitions and Basic Properties
Definition 2.1.
The Lasso estimator is noted as $\hat{\beta}_{\mathrm{L}}$, with its definition as follows:
$\hat{\beta}_{\mathrm{L}}=\underset{\beta\in\mathbb{R}^{p}}{\arg\min}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\lambda_{\mathrm{L}}\|\beta\|_{1}\right\},$
Proposition 2.2.
The KKT conditions for the Lasso estimator follow from subdifferential calculus (van de Geer, 2016). They can be expressed as
$\frac{1}{n}X^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}\bar{z},$
where $\bar{z}\in\partial\|\hat{\beta}_{\mathrm{L}}\|_{1}$, with the $i$-th component given by
$\bar{z}_{i}=\begin{cases}\mathrm{sign}((\hat{\beta}_{\mathrm{L}})_{i}),&\text{if }(\hat{\beta}_{\mathrm{L}})_{i}\neq 0,\\[6.0pt] u_{i}\in[-1,1],&\text{if }(\hat{\beta}_{\mathrm{L}})_{i}=0,\end{cases}\qquad\text{for }i=1,\dots,p.$
Also, notice that:
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Here, $\mathrm{sign}(\cdot)$ denotes the sign function. In addition, the KKT condition on the active block derives
$\hat{\beta}_{\mathrm{L},E}=(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s).$
Assumption 2.3.
The columns of the design matrix $X$ are in general position.
Proposition 2.4 (Uniqueness of the Lasso).
Under Assumption 2.3, the Lasso solution $\hat{\beta}_{\mathrm{L}}$ is unique, with support exactly $E$; that is, $\hat{\beta}_{j}\neq 0$ iff $j\in E$. Also, $\mathrm{null}(X_{E})=\{0\}$. See Tibshirani (2013) for further details.
Remark 2.5.
Assumption 2.3 ensures that the definition of the refined estimator is unambiguous. Throughout this paper, we assume that Assumption 2.3 holds without further mention.
Definition 2.6.
We define the key quantities associated with the Lasso:
  • The equicorrelation set:
    $E\coloneqq\{\,i\in\{1,2,\ldots,p\}:|\frac{1}{n}X_{i}^{\top}(y-X\hat{\beta}_{\mathrm{L}})|=\lambda_{\mathrm{L}}\,\}.$
  • Restricted Gram matrix:
    $\Sigma_{n,E}\coloneqq\frac{X_{E}^{\top}X_{E}}{n}.$
  • The equicorrelation signs:
    $s\coloneqq\operatorname{sign}({X_{E}}^{T}(y-X\hat{\beta}_{\mathrm{L}})).$
Proposition 2.7.
By the polyhedral lemma (Lee et al., 2016), conditioning on a fixed active set $E=E_{i}$ and sign vector $s=s_{i}$ is equivalent to restricting the random noise vector $\varepsilon$ to lie in a corresponding polyhedral region. Specifically, there exist a matrix $A_{i}$ and a vector $b_{i}$ such that
$\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}=\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}=\mathbf{1}_{\{\varepsilon\in\mathcal{P}_{i}\}},$
where $\mathcal{P}_{i}=\{\varepsilon:A_{i}\varepsilon\leq b_{i}\}$ denotes the polyhedron associated with the selection event $\{E=E_{i},s=s_{i}\}$.
Proposition 2.8.
The ridge regression estimator (Hoerl and Kennard, 1970) is defined as
$\hat{\beta}^{\mathrm{r}}=\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda^{\mathrm{r}}}{2}\|\beta\|_{2}^{2}\right\},$
where $\lambda^{\mathrm{r}}>0$ is a regularization parameter. The corresponding Karush–Kuhn–Tucker (KKT) condition is
$-\frac{1}{n}X^{\top}\bigl(y-X\hat{\beta}^{\mathrm{r}}\bigr)+\lambda^{\mathrm{r}}\hat{\beta}^{\mathrm{r}}=0,$
which yields the closed-form expression
$\hat{\beta}^{\mathrm{r}}=(X^{\top}X+n\lambda^{\mathrm{r}}I_{p})^{-1}(X^{\top}y).$
2.3 The Lasso–Ridge Estimator and Related Definitions
Definition 2.9.
A Lasso-Ridge estimator could be defined as:
$\hat{\beta}_{\mathrm{R}}=\underset{\beta\in\mathbb{R}^{p},\beta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\beta-\hat{\beta}_{\mathrm{L}}\|_{2}^{2}\right\},$
or equivalently:
$\hat{\beta}_{\mathrm{R}}=\hat{\delta}+\hat{\beta}_{\mathrm{L}},$
$\quad\hat{\delta}=\underset{\delta\in\mathbb{R}^{p},\delta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-X\hat{\beta}_{\mathrm{L}}-X\delta\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta\|_{2}^{2}\right\}.$
Definition 2.10.
Let $\hat{\beta}_{\mathrm{L}}$ and $\hat{\beta}_{\mathrm{R}}$ Lasso and Lasso–Ridge estimators of $\beta_{0}$. We define the prediction error gap:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}}):=\frac{1}{n}(\|X(\hat{\beta}_{\mathrm{L}}-\beta_{0})\|_{2}^{2}-\|X(\hat{\beta}_{\mathrm{R}}-\beta_{0})\|_{2}^{2}).$
Define the difference in prediction mean squared error:
$\mathrm{DMSE}:=\frac{1}{n}\,\mathbb{E}[\|X(\hat{\beta}_{\mathrm{L}}-\beta_{0})\|_{2}^{2}-\|X(\hat{\beta}_{\mathrm{R}}-\beta_{0})\|_{2}^{2}].$
A positive $\mathrm{DMSE}$ means $\hat{\beta}_{\mathrm{R}}$ strictly improves prediction MSE in expectation.
Definition 2.11 (Projection operators).
Let $A\in\mathbb{R}^{n\times k}$ be a matrix with full column rank, i.e., $\operatorname{rank}(A)=k$. The orthogonal projection onto the column space of $A$ is defined by
$P_{A}:=A(A^{\top}A)^{-1}A^{\top}.$
The projection onto the orthogonal complement of $\operatorname{col}(A)$ is
$P_{A}^{\perp}:=I_{n}-P_{A}.$
For any $u\in\mathbb{R}^{n}$, $P_{A}u$ is the projection of $u$ onto $\operatorname{col}(A)$, and $P_{A}^{\perp}u$ is the corresponding orthogonal residual. In particular, if $X_{E}\in\mathbb{R}^{n\times|E|}$ denotes the submatrix of $X$ with columns indexed by $E$, and $\operatorname{rank}(X_{E})=|E|$, then
$P_{X_{E}}$ $=X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top},$
$P_{X_{E}}^{\perp}$ $=I_{n}-P_{X_{E}}.$
3 Main Results
Theorem 3.1.
Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have
$\mathrm{DMSE}$ $\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})})\Bigr\}.$
Proof.
We start by rewriting the target term as follows:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$ $=\frac{1}{n}\|X\beta_{0}-X\hat{\beta}_{\mathrm{L}}\|_{2}^{2}-\frac{1}{n}\|X\beta_{0}-X\hat{\beta}_{\mathrm{L}}-X\hat{\delta}\|_{2}^{2}$
$=\frac{2}{n}\langle X\beta_{0}-X\hat{\beta}_{\mathrm{L}},X\hat{\delta}\rangle-\frac{1}{n}\langle X\hat{\delta},X\hat{\delta}\rangle$
$=\frac{2}{n}\langle X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E},X_{E}\hat{\delta}_{E}\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$ (1)
Recalling the closed-form expression for the nonzero part of the Lasso estimator under the KKT conditions (Proposition 2.2), we have
$\hat{\beta}_{\mathrm{L},E}=(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s).$
The term $X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E}$ appearing in the previous step can be rewritten as follows:
$X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E}$ $=X\beta_{0}-X_{E}(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s)$
$=X\beta_{0}-X_{E}(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}(X\beta_{0}+\varepsilon)-n\lambda_{\mathrm{L}}s)$
$=P_{X_{E}}^{\perp}X\beta_{0}-P_{X_{E}}\varepsilon+n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,$
where $P_{X_{E}}$ denotes the orthogonal projection onto $\operatorname{col}(X_{E})$ and $P_{X_{E}}^{\perp}=I_{|E|}-P_{X_{E}}$ is the projection onto its orthogonal complement. Inserting this rewritten form into equality (1) and rearranging, we obtain
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$ $=\frac{2}{n}\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle-\frac{2}{n}\langle P_{X_{E}}\varepsilon,X_{E}\hat{\delta}_{E}\rangle$
$+\frac{2}{n}\langle n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,X_{E}\hat{\delta}_{E}\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$ (2)
Next, we deal with the four inner products appearing in the previous equation (2). First, note that for the first term, we have
$\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(I_{|E|}-X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
$=\hat{\delta}_{E}^{\top}(X_{E}^{\top}-X_{E}^{\top}X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
$=\hat{\delta}_{E}\mathbf{0}_{|E|}X\beta_{0}$
$=0.$
Next, we simplify the second term in equation (2) using straightforward matrix multiplication:
$\langle P_{X_{E}}\varepsilon,X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})\varepsilon$
$=\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
We now simplify the third term:
$\frac{2}{n}\langle n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,X_{E}\hat{\delta}_{E}\rangle$ $=2\lambda_{\mathrm{L}}\hat{\delta}_{E}^{\top}X_{E}^{\top}X_{E}(X_{E}^{\top}X_{E})^{-1}s$
$=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle.$
Substituting the above three equalities into equation (2), we obtain
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle-\frac{2}{n}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$ (3)
Consider the case where $E\neq\emptyset$. Then $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ admits the orthogonal decomposition:
$\Sigma_{n,E}=P\operatorname{diag}(\mu_{1}+,\dots,\mu_{|E|})P^{\top}$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the is the corresponding eigenvector matrix. According to Lemma A.3 in the Appendix, we obtain the following closed-form representation of $\hat{\delta}_{E}$:
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
By orthogonal decomposition, we have
$2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle$ $=2\lambda_{\mathrm{L}}^{2}s^{\top}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=2\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{1}{\mu_{1}+\lambda_{\mathrm{R}}},\dots,\frac{1}{\mu_{|E|}+\lambda_{\mathrm{R}}})P^{\top}s.$
Following a similar argument, we derive
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$ $=\frac{1}{n}\hat{\delta}_{E}^{\top}X_{E}^{\top}X_{E}\hat{\delta}_{E}$
$=\lambda_{\mathrm{L}}^{2}s^{\top}(\lambda_{\mathrm{R}}\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}\Sigma_{n,E}(\lambda_{\mathrm{R}}\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s.$
For notational convenience, define
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}}):=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$
Combining it with the previous two equations, we obtain the following simplified expression:
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}+2\lambda_{\mathrm{R}}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}+2\lambda_{\mathrm{R}}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s.$
Consider the function
$f_{1}(x)=\frac{x+2a}{(x+a)^{2}}.$
Notice that $\mu_{1}+\cdots+\mu_{|E|}=\operatorname{trace}(\Sigma_{n,E})=|E|$. Since $f_{1}(x)$ is monotone decreasing, for any $\mu_{1},\dots,\mu_{|E|}\leq|E|$,
$f_{1}(\mu_{i})\geq f_{1}(|E|),\quad i=1,\dots,|E|.$
Recall that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ where $c>2$. Under this choice, we derive the following inequalities:
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}+2c|E|}{(\mu_{1}+c|E|)^{2}},\dots,\frac{\mu_{|E|}+2c|E|}{(\mu_{|E|}+c|E|)^{2}})P^{\top}s$
$\geq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{|E|+2c|E|}{(|E|+c|E|)^{2}},\dots,\frac{|E|+2c|E|}{(|E|+c|E|)^{2}})P^{\top}s$
$=\frac{\lambda_{\mathrm{L}}^{2}}{|E|}\frac{2c+1}{(c+1)^{2}}\|P^{\top}s\|_{2}^{2}$
$=\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}.$
Recall that the above inequality holds on the event $\{E\neq\emptyset\}$. Since $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})=0$ on $\{E=\emptyset\}$, we have
$\mathbb{E}[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})]$ $=\mathbb{E}\big[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E\neq\emptyset\}}+H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E=\emptyset\}}\big]$
$=\mathbb{E}[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E\neq\emptyset\}}]$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\,\mathbb{P}(E\neq\emptyset).$ (4)
Next, we consider the remaining part of equation (3). From Lemma A.3 in the appendix, we know the following equation holds when $\lambda_{\mathrm{R}}>2|E|$:
$\|\hat{\delta}_{E}\|_{1}=s^{\top}\hat{\delta}_{E}.$
Since $\|\hat{\delta}_{E}\|_{1}=0$ on $\{E=\emptyset\}$, it is enough to consider the case $E\neq\emptyset$ for the upper bound:
$\|\hat{\delta}_{E}\|_{1}$ $=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}s^{\top}P\operatorname{diag}(\frac{\lambda_{\mathrm{R}}}{\mu_{1}+\lambda_{\mathrm{R}}},\dots,\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}})P^{\top}s$
$\leq\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}}\|P^{\top}s\|_{2}^{2}$
$\leq\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}}|E|$
$\leq\frac{\lambda_{\mathrm{L}}}{c}.$
It follows from the dual-norm inequality that
$\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\mathbb{E}\langle\hat{\delta}_{E},X_{E}^{\top}\varepsilon\rangle$
$\leq\mathbb{E}[\|\hat{\delta}_{E}\|_{1}\,\|X_{E}^{\top}\varepsilon\|_{\infty}]$
$\leq\frac{\lambda_{\mathrm{L}}}{c}\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}].$ (5)
Next, we analysis the term $\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ in the following way:
$\|X_{E}^{\top}\varepsilon\|_{\infty}$ $=\|X_{E}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\subseteq T_{0}\}}+\|X_{E}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}$
$\leq\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\subseteq T_{0}\}}+\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}$
$\leq\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}+\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}.$
By taking expectations and applying the Cauchy–Schwarz inequality, we obtain
$\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ $\leq\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}]$
$\leq\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\not\subseteq T_{0}\}}^{2}]}$
$=\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}.$
Since the inequality holds for every subset $T_{0}\subseteq\{1,\dots,p\}$, we may take the infimum over all such $T_{0}$ and derive
$\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ $\leq\inf_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr\}.$ (6)
Combining $(3)$, $(4)$, $(5)$, and $(6)$, we obtain the following inequalities, which conclude the proof:
$\mathrm{DMSE}$ $=\mathbb{E}\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$
$=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\,\mathbb{P}(E\neq\emptyset)-\frac{2\lambda_{\mathrm{L}}}{cn}\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$
$\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})})\Bigr\}.$
Lemma 3.2.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ and $\|X_{j}\|_{2}^{2}=n$ for all $j$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$ See Lemma A.1 for a proof.
Remark 3.3 (Approximate support containment).
Let $S_{0}=\{j:\beta_{0,j}\neq 0\}$ denote the true support. We assume the existence of a deterministic index set $T_{0}$ such that
$\mathbb{P}(E\subseteq T_{0})\geq 1-\delta$
for a fixed error probability $0<\delta<1$. When $T_{0}=S_{0}$, $E\subseteq S_{0}$ corresponds to the event that the Lasso has no false alarms, which occurs with high probability under mild conditions (van de Geer, 2016, Section 4.7, p. 51). Next, considering the case where $T_{0}=\{1,\dots,p\}$, we obtain the following corollary:
Corollary 3.4.
Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have:
$\mathrm{DMSE}$ $\geq\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X^{\!\top}\varepsilon\|_{\infty}]).$
Corollary 3.5.
Suppose the penalty level $\lambda_{\mathrm{L}}$ of the Lasso estimator satisfies
$\lambda_{\mathrm{L}}>\frac{\mathbb{E}[\|X^{\!\top}\varepsilon\|_{\infty}]}{n\mathbb{P}(E\neq\emptyset)}.$
Then there exists a constant $c$ for the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ such that
$\mathrm{DMSE}>0.$
In particular, the Lasso cannot achieve the optimal prediction risk whenever $\lambda_{\mathrm{L}}$ slightly exceeds the the expectation of the random maximum.
Remark 3.6.
For further discussions, we first recall the standard bound for $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]$. We then discuss how the structure of the design matrix $X$ critically influences the behavior of this Gaussian maximum.
Lemma 3.7 (Expectation of Gaussian maxima).
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $X=[x_{1},\ldots,x_{p}]\in\mathbb{R}^{n\times p}$ with column norms $\|x_{j}\|_{2}\leq\sqrt{n}$ for all $j$. Then
$\frac{1}{n}\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]\leq\sigma\sqrt{\frac{2\log(2p)}{n}}.$
For more details, see (van de Geer, 2016, Section 17.4, p. 243).
Lemma 3.8 (Generic chaining bound).
Let $(T,d)$ be a metric space. If a centered process $\{X_{t}:t\in T\}$ has sub-Gaussian increments with respect to $d$, then there exists an absolute constant $C>0$ such that
$\mathbb{E}\sup_{t\in T}X_{t}\;\leq\;C\,\gamma_{2}(T,d),$
where $\gamma_{2}(T,d)$ is the generic chaining functional
$\gamma_{2}(T,d)=\inf_{\{T_{k}\}}\sup_{t\in T}\sum_{k\geq 0}2^{k/2}\,d(t,T_{k}),$
the infimum being over admissible sequences $\{T_{k}\}$, that is, $|T_{0}|=1$ and $|T_{k}|\leq 2^{2^{k}}$. For more details, see (Vershynin, 2018, Theorem 8.5.3, p. 208).
Lemma 3.9 (Two-step bound under column normalization and clustering).
Let $X=[X_{1},\dots,X_{p}]\in\mathbb{R}^{n\times p}$ satisfy $\|X_{j}\|_{2}=\sqrt{n}$ for all $j$. Assume that the index set $\{1,\dots,p\}$ is partitioned into $K$ clusters $\{\mathcal{C}_{\ell}\}_{\ell=1}^{K}$, and for each cluster fix a representative $r_{\ell}\in\mathcal{C}_{\ell}$. Write $r(j)$ for the representative of the cluster containing $j$, and suppose the clusters are $\delta$-tight in the sense that
$\|X_{j}-X_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j,$
for some $\delta\in(0,1]$. Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Then
$\frac{1}{n}\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]\leq\frac{\sigma}{\sqrt{n}}(\sqrt{2\log(2K)}+\delta\sqrt{2\log(2p)}).$
Remark 3.10.
The preceding lemmas indicate that the structure of the design matrix affects the choice of the Lasso tuning parameter. In particular, when the columns of the design matrix are more strongly correlated, a smaller tuning parameter is required, since overly large tuning values may lead to inadmissibility. In addition, a direct calculation of the prediction improvement is possible; see Lemma A.5 for further details. Finally, we consider a special case in which the Lasso selects the true support with high probability, taking into account the symmetry of the noise in addition to its tail behavior.
Theorem 3.11.
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, then we have
$\mathrm{DMSE}$ $\geq\frac{\lambda_{\mathrm{L}}}{\sqrt{c}}(\frac{\sqrt{c}(2c+1)}{(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\sigma(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}).$
Remark 3.12.
This theorem shows that when the Lasso selects the true support with high probability, its prediction performance can be improved by a suitable refinement over a wide range of tuning parameters. Moreover, it is known that variable selection consistency may occur even when the Lasso tuning parameter does not vanish; see, for example, (Lahiri, 2021). In such settings, the present theorem implies that prediction improvement can be achieved.
4 Extensions
The analysis developed in this paper is not limited to the classical Lasso estimator. Its core arguments extend naturally to other $\ell_{1}$-based procedures, most notably the Dantzig selector and the Adaptive Lasso. We describe these extensions below.
Dantzig selector.
The Dantzig selector (Candès and Tao, 2007) is defined by
$\hat{\beta}_{\mathrm{D}}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{p}}\|\beta\|_{1}\quad\text{subject to}\quad\left\|X^{\top}(y-X\beta)/n\right\|_{\infty}\leq\lambda_{\mathrm{D}}.$
With the equalcorrelation set defined analogously under the Dantzig tuning parameter $\lambda_{\mathrm{D}}$, all arguments and results developed above continue to hold with only notational modifications.
Adaptive Lasso.
The Adaptive Lasso (Zou, 2006) solves
$\hat{\beta}_{\mathrm{A}}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{p}}\Bigl\{\|y-X\beta\|_{n}^{2}+\lambda_{\mathrm{A}}\sum_{j=1}^{p}w_{j}|\beta_{j}|\Bigr\},$
where $w_{j}>0$ are data-dependent weights, typically derived from an initial estimator. For columns of $X$ corresponding to a single tuning weight, or more generally for groups of columns associated with the same effective penalty level, an equalcorrelation set can be defined in a manner analogous to the unweighted case. As a result, the framework of Theorem 3.1 and the main analysis extend directly, and the corresponding DMSE calculations remain valid with only notational modifications.
5 Conclusions
This paper revisits the classical tuning paradigm for the Lasso through direct comparisons of prediction risk. We show that, in certain regimes, the Lasso equipped with the conventional choice $\lambda\asymp\sigma\sqrt{\log(p)/n}$ can be dominated in mean squared prediction error. Our analysis is based on an explicit DMSE decomposition, which identifies settings in which stronger regularization leads to suboptimal prediction performance. From this viewpoint, these phenomena provide a quantitative explanation for empirical observations that smaller tuning parameters may perform favorably in correlated designs and some other settings.
Although the Lasso serves as a primary reference, the arguments developed here are not specific to a single estimator and point to a broader perspective on $\ell_{1}$-regularized methods. Extensions to more general settings, including models with dependent or heteroscedastic noise and nonlinear frameworks, remain an interesting direction for future work.
Appendix
Appendix A Proofs of Technical Lemmas
Lemma A.1.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ and $\|X_{j}\|_{2}^{2}=n$ for all $j$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$
Proof.
For each $j$,
$\mathbb{P}(|X_{j}^{\top}\varepsilon|\geq t)$ $\leq 2\exp(-t^{2}/(2\sigma^{2}n)).$
Define $M=\|X^{\top}\varepsilon\|_{\infty}$. Then
$\mathbb{P}(M\geq t)$ $\leq 2p\exp(-t^{2}/(2\sigma^{2}n)).$
Using $\mathbb{E}[M^{2}]=\int_{0}^{\infty}2t\,\mathbb{P}(M\geq t)\,dt$, let $t_{0}=\sigma\sqrt{n}\sqrt{2\log(4p)}$. Then
$\int_{0}^{t_{0}}2t\,\mathbb{P}(M\geq t)\,dt$ $\leq t_{0}^{2}=2\sigma^{2}n\log(4p),$
$\int_{t_{0}}^{\infty}2t\,\mathbb{P}(M\geq t)\,dt$ $\leq 4p\int_{t_{0}}^{\infty}t\exp(-t^{2}/(2\sigma^{2}n))\,dt$
$=4p\,\sigma^{2}n\int_{\log(4p)}^{\infty}e^{-u}\,du$
$=\sigma^{2}n.$
Combining these two inequalities, we have
$\mathbb{E}[M^{2}]$ $\leq\sigma^{2}n\,(2\log(4p)+1),$
which concludes the proof. ∎
Lemma A.2 (Newmann Serious Expansion).
Let $A\in\mathbb{R}^{d\times d}$ with $\|A\|_{\infty}<1$. Then we have
$(I-A)^{-1}=\sum_{k=0}^{\infty}A^{k}.$
For more details, see (Gentle, 2024, Section 3.11.6, p. 197).
Lemma A.3 (Closed-form Representation under the KKT Conditions).
When $E\neq\emptyset$, the nonzero entries of $\hat{\delta}$ admit a simple expression:
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
Proof.
Following from the second definition of our refitting step, we have
$\hat{\delta}=\underset{\delta\in\mathbb{R}^{p},\delta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-\hat{\beta}_{\mathrm{L}}-X_{E}\delta_{E}\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta_{E}\|_{2}^{2}\right\},$
So when $E\neq\emptyset$, $\hat{\delta}_{E}$ could be regarded as a simple ridge regression estimator:
$\hat{\delta}_{E}=\underset{\delta\in\mathbb{R}^{|E|}}{\arg\min}\left\{\frac{1}{2n}\|y-X\hat{\beta}_{\mathrm{L}}-X_{E}\delta_{E}\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta_{E}\|_{2}^{2}\right\}.$
Recall the closed from representation under the KKT conditions of a ridge regression model in Proposition 2.8, we derive
$\hat{\delta}_{E}=\frac{1}{n}(\frac{1}{n}X_{E}^{\top}X_{E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}}).$
From the KKT conditions of the Lasso step, we obtain the following result:
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s,$
Combining the previous two equations. Also replacing $X_{E}^{\top}X_{E}/n$ with $\Sigma_{n,E}$ and factoring $\lambda_{\mathrm{R}}$ out of the inverse completes the proof. ∎
Lemma A.4 (Alternative Form of $\|\hat{\delta}_{E}\|_{1}$).
Consider the case where $E\neq\emptyset$ and $\lambda_{\mathrm{R}}>2|E|$. We have $\|\hat{\delta}_{E}\|_{1}=s^{\top}\hat{\delta}_{E}$.
Proof.
Notice that, it suffices to show that $\operatorname{sign}(s)=\operatorname{sign}(\hat{\delta}_{E})$. According to lemma A.3, we have:
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{R}}}{\lambda_{\mathrm{L}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
Thus, the problem can be reformulated as follows:
$\operatorname{sign}((\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s)=\operatorname{sign}(s).$
Notice that $\|\Sigma_{n,E}\|_{\infty}\leq|E|$, since $\|\Sigma_{n,E}\|_{\infty}=\max_{i\in E}\sum_{j\in E}|(\Sigma_{n,E})_{ij}|$ and each $|(\Sigma_{n,E})_{ij}|\leq 1$. When $\lambda_{\mathrm{R}}>2|E|$, we have
$\|\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}\|_{\infty}<\frac{1}{2}.$
Applying Newmann series expansion (Lemma A.2), we get:
$(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}=\sum_{k=0}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}=I_{|E|}+\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}.$
Thus, we obtain
$(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s=s+(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s,$
and our target is reduced to showing that
$\operatorname{sign}(s+(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s)=\operatorname{sign}(s).$
If each entry of the second term on the left-hand side has absolute value less than $1$, then the preceding equality follows. Hence, it suffices to prove that
$\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s\|_{\infty}<1.$
Noting that $s$ is a binary vector, we have
$\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s\|_{\infty}$ $\leq\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})\|_{\infty}\leq\sum_{k=1}^{\infty}\|(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}\|_{\infty}$
$\leq\sum_{k=1}^{\infty}\|(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})\|_{\infty}^{k}<\sum_{k=1}^{\infty}(\frac{1}{2})^{k}=1.$
Lemma A.5 (Expectation of a Gaussian linear form under polyhedral truncation).
Recall that we have the following definition:
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}}):=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$
We have derived the following result in equation (3) of the main text:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})=H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
Notice that both $E_{i}$ and $s_{i}$ are determined by $\lambda_{\mathrm{L}}$, and this dependence is implicit in the notation $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{D}})$. Moreover, on the event where $E_{i}$ and $s_{i}$ remain constant, this term is also determined as a constant. Recall the polyhedral lemma (Lee et al., 2016), let $\{(E_{i},s_{i})\}_{i=1}^{3^{p}}$ be an enumeration of all possible support–sign pairs, we have
$\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}=\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}=\mathbf{1}_{\{\varepsilon\in\mathcal{P}_{i}\}},$
where $\mathcal{P}_{i}=\{\varepsilon:A_{i}\varepsilon\leq b_{i}\}$ denotes the polyhedron associated with the selection event $\{E=E_{i},s=s_{i}\}$. Collecting the above definitions and results, we obtain
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\sum_{i}\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\sum_{i}\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}$
$=\sum_{i}\mathbb{E}H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\sum_{i}\mathbb{E}\langle X_{E_{i}}\hat{\delta}_{E_{i}},\varepsilon\mathbf{1}_{\{E=E_{i},\,s=s_{i}\}}\rangle$
$=\sum_{i}\mathbb{E}H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\sum_{i}\mathbb{E}\langle X_{E_{i}}\hat{\delta}_{E_{i}},\varepsilon\mathbf{1}_{\{A_{i}\varepsilon\leq b_{i}\}}\rangle,$
where $H_{E_{i},s_{i}}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ denotes the value of $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ on the corresponding event.
Appendix B Proofs of the Main Theorems
Proof of corollary 3.5
Consider the function
$f_{2}(x)=\frac{x(2x+1)}{(x+1)^{2}},\quad x\geq 0.$
The function $f_{2}(x)$ is monotone increasing, and it converges to $2$ as $x\to\infty$. Taking this back to Theorem 2.2 concludes the proof.
Proof of Lemma 3.9
Proof.
Recall that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and define, for each $j\in\{1,\dots,p\}$,
$Z_{j}:=\frac{1}{n}X_{j}^{\top}\varepsilon,$
where $X_{j}\in\mathbb{R}^{n}$ denotes the $j$-th column of $X$. For each $j$, decompose
$|Z_{j}|=\big|Z_{r(j)}+(Z_{j}-Z_{r(j)})\big|\leq|Z_{r(j)}|+|Z_{j}-Z_{r(j)}|.$
Taking maxima and expectations,
$\mathbb{E}\max_{j}|Z_{j}|\leq\mathbb{E}\max_{\ell\leq K}|Z_{r_{\ell}}|+\mathbb{E}\max_{j\leq p}|Z_{j}-Z_{r(j)}|.$
First, note that each $Z_{r_{\ell}}$ is centered Gaussian with variance $\sigma^{2}/n$. By Lemma 3.7 with $p$ replaced by $K$,
$\mathbb{E}\max_{\ell\leq K}|Z_{r_{\ell}}|\leq\frac{\sigma}{\sqrt{n}}\sqrt{2\log(2K)}.$
Second, for each $j$,
$Z_{j}-Z_{r(j)}=\frac{1}{n}(X_{j}-X_{r(j)})^{\top}\varepsilon,\quad\operatorname{Var}(Z_{j}-Z_{r(j)})=\frac{\sigma^{2}}{n^{2}}\|X_{j}-X_{r(j)}\|_{2}^{2}\leq\frac{\sigma^{2}}{n}\delta^{2}.$
Thus $\{Z_{j}-Z_{r(j)}\}_{j=1}^{p}$ are centered Gaussian variables with variance at most $\sigma^{2}\delta^{2}/n$. By Lemma 3.7,
$\mathbb{E}\max_{j\leq p}|Z_{j}-Z_{r(j)}|\leq\frac{\sigma}{\sqrt{n}}\,\delta\,\sqrt{2\log(2p)}.$
Combining the two bounds completes the proof:
$\mathbb{E}\max_{1\leq j\leq p}|Z_{j}|\leq\frac{\sigma}{\sqrt{n}}\big(\sqrt{2\log(2K)}+\delta\sqrt{2\log(2p)}\big).$
Proof of theorem 2.11
We first recall the following expression for the DMSE:
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\mathbb{P}(E\neq\emptyset)-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$ (1)
Also, recall that according to Lemma A.3, we have the closed-form expression for the Lasso–Ridge estimator:
$\hat{\delta}_{E}=\lambda_{\mathrm{L}}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s,$
which is determined once the values of $s$ and $|E|$ are fixed. We consider an alternative approach for the following term, with the tuning parameter $\lambda_{\mathrm{R}}$ in the second step set to $c|E|$:
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]+\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
$=\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]+\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}],$
where and we adopt the same notation for any index set, in particular, $\Sigma_{n,S_{0}}\coloneqq X_{S_{0}}^{\top}X_{S_{0}}/n$, and $s_{0}$ denotes the sign vector corresponding to $s$ on the event $E={S}_{0}$. Note that, by symmetry of the Gaussian distribution, we have
$\frac{1}{n}\mathbb{E}\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle=0,$
Substituting this into the previous equation, we obtain
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]-\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}].$
First, we prove an upper bound of $\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$. Since $\|X_{E}\hat{\delta}_{E}\|_{2}^{2}=0$ on $\{E=\emptyset\}$, it is enough to consider the case $E\neq\emptyset$, where we have the following orthogonal decomposition based on the closed-form expression of the Lasso–Ridge estimator in Lemma A.3:
$\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$ $=\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$
$=\lambda_{\mathrm{L}}^{2}s^{\top}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}\Sigma_{n,E}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s.$
Consider the function
$f_{3}(x)=\frac{x}{(x+a)^{2}},\quad a>0.$
We observe that $f_{3}(x)$ attains its maximum at $f(a)$. Thus, we have
$\frac{\mu_{i}}{(\mu_{i}+c|E|)^{2}}\leq\frac{c|E|}{(c|E|+c|E|)^{2}},\quad i=1,\dots,|E|.$
Based on this, and recalling that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$, we derive
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+c|E|)^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+c|E|)^{2}})P^{\top}s$
$\leq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{c|E|}{(c|E|+c|E|)^{2}},\dots,\frac{c|E|}{(c|E|+c|E|)^{2}})P^{\top}s$
$\leq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{c|E|}{(2c|E|)^{2}},\dots,\frac{c|E|}{(2c|E|)^{2}})P^{\top}s$
$\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c|E|}\|P^{\top}s\|_{2}^{2}$
$\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c}.$ (3)
Also note that, since $\|\varepsilon\|_{2}^{2}/\sigma^{2}\sim\chi^{2}_{n}$, we have
$\mathbb{E}\!\left[\|\varepsilon\|_{2}^{2}\right]=n\sigma^{2}.$
Now returning to our target, we have
$\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$ $\leq\frac{1}{n}\sqrt{\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\neq S_{0}\}}^{2}]}$
$=\frac{1}{n}\sqrt{\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{1}{n}\sqrt{\mathbb{E}[\|X_{E}\hat{\delta}_{E}\|_{2}^{2}\,\|\varepsilon\|_{2}^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\sqrt{\frac{\lambda_{\mathrm{L}}^{2}}{4c}\mathbb{E}[\frac{1}{n}\|\varepsilon\|_{2}^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}\sqrt{\mathbb{P}(E\neq S_{0})}.$ (4)
Next, we derive an upper bound for
$\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
using a similar approach. First, following an argument similar to that used to derive inequality (3), we obtain
$\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c}.$
Since $\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle$ is a Gaussian random variable, we have
$\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle^{2}]$ $=\sigma^{2}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}.$
Thus for the second term of our final target, we derive
$\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]$
$\leq\frac{1}{n}\sqrt{\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\neq S_{0}\}}^{2}]}$
$\leq\frac{1}{n}\sqrt{\sigma^{2}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma}{\sqrt{n}}\sqrt{\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{cn}}\sqrt{\mathbb{P}(E\neq S_{0})}.$ (5)
Combining (4) and (5), we have
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]-\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}\sqrt{\mathbb{P}(E\neq S_{0})}+\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{cn}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}.$
Substituting this into equation (1), we obtain
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\mathbb{P}(E\neq\emptyset)-\frac{\sigma\lambda_{\mathrm{L}}}{\sqrt{c}}(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}$
$=\frac{\lambda_{\mathrm{L}}}{\sqrt{c}}(\frac{\sqrt{c}(2c+1)}{(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\sigma(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}).$
Consider the function
$f_{3}(x)=\frac{\sqrt{x}(2x+1)}{(x+1)^{2}},\qquad x\geq 0.$
Note that $f_{3}(x)$ increases before $x\approx 0.79$, decreases after $x\approx 0.79$, and attains its maximum value $f_{3,\max}\approx 1.78$ at $x\approx 0.79$.