### v2

# On the Prediction Suboptimality of the Lasso

Abstract
The choice of the tuning parameter in the Lasso is central to its statistical performance in high-dimensional linear regression. Classical consistency theory characterizes the rate of the Lasso tuning parameter, and numerous works provide non-asymptotic guarantees. However, the optimal choice of tuning within a fully non-asymptotic framework remains incompletely understood. In this work, we study tuning regimes under which the Lasso exhibits suboptimal prediction performance, in the sense that it is dominated in mean squared prediction error by a simple refinement. We further examine how structural factors in the design matrix influence the suboptimality phenomenon, and discuss extensions to other estimators and more general noise structures.
1 Introduction
Linear regression is a simple yet powerful framework for high-dimensional modeling, valued for its interpretability and its transparent representation of signal and noise. The Lasso (Tibshirani, 1996) is one of the most widely used methods for high-dimensional linear regression. By combining estimation and variable selection through an $\ell_{1}$ penalty, it yields computationally efficient estimators with well-developed theoretical guarantees. A large body of work has established oracle inequalities, prediction error bounds, and sparsity recovery results under various assumptions on the design matrix; see, for example, (van de Geer, 2016) and the references therein. Within this literature, it is common to select the tuning parameter at the universal scale $\sigma\sqrt{(\log p)/n}$, which ensures favorable asymptotic and non-asymptotic guarantees in high-dimensional settings.
More recently, it has been observed that the choice of the tuning parameter can be more delicate, particularly when the design matrix exhibits strong correlations. For instance, related results in (van de Geer and Lederer, 2013; Hebiri and Lederer, 2013) indicate that stronger correlations in the design call for smaller tuning parameters in order to obtain tighter guarantees. In a related direction, (Dalalyan et al., 2017) consider new measures of design correlation and establish refined Lasso oracle inequalities within this framework. From a different perspective, (Bellec et al., 2018) show that adaptively chosen tuning parameters for the Lasso can yield sharp oracle inequalities and improved minimax prediction rates under suitable design conditions. Taken together, these results suggest that tuning parameters below the universal Lasso scale of order $\sigma\sqrt{(\log p)/n}$ may be advantageous in certain regimes. These works primarily focus on refined risk upper bounds and therefore do not fully address this aspect of the problem, and we are interested in whether the presence of $\ell_{1}$-regularization inherently induces a broad range of tuning parameters that lead to suboptimal prediction.
In this paper, we provide new insights into the prediction and tuning of $\ell_{1}$-regularized high-dimensional regression from the perspective of prediction risk dominance within a non-asymptotic framework. Rather than proposing new tuning rules for refined upper bounds, we characterize the range of tuning parameters for which prediction suboptimality occurs and establish lower bounds on the prediction mean squared error difference. Specifically, for the Lasso, when the tuning parameter exceeds approximately $\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}/n],$ where $T_{0}$ denotes a high-probability containment set of the selected support of the Lasso, it can be dominated in prediction mean squared error by a suitably constructed refinement, as suggested in (Liu, 2025). Notably, the lower bound is explicit, with no hidden constants, and the resulting dominance range can include tuning levels at the standard theoretical scale $\lambda\asymp\sigma\sqrt{(\log p)/n}$.
Although the Lasso serves as the primary example, the analysis may extend to a broader class of $\ell_{1}$-regularized estimators and to more general noise structures. The results provide a theoretical explanation for the empirical observation that smaller tuning parameters can yield improved prediction performance, while clarifying the distinction between rate-optimality and finite-sample risk dominance. The main contributions are as follows:
  1. 1.
    A framework for direct comparison of prediction risks, allowing one to assess whether a given Lasso tuning choice admits improvement via a simple refinement.
  2. 2.
    Explicit range conditions on the tuning parameter under which the Lasso is dominated in mean squared prediction error.
  3. 3.
    A characterization of how structural features, including sparsity and design correlations, shape the regimes of prediction risk dominance, together with several extensions.
2 Preliminaries and Background
2.1 Notation and Model Assumptions
The cardinality of a set $E$ is denoted by $|E|$. For a vector $x$, define $\|x\|_{1}=\sum_{i}|x_{i}|$ and $\|x\|_{2}=\bigl(\sum_{i}x_{i}^{2}\bigr)^{1/2}$. For a matrix $A=(A_{ij})\in\mathbb{R}^{n\times p}$, define
$\|A\|_{\infty}$ $=\max_{i}\sum_{j}|A_{ij}|,$
$\|A\|_{2}$ $=\sup_{\|x\|_{2}=1}\,\|Ax\|_{2}.$
For any $S\subseteq[p]$, $A_{S}$ denotes the submatrix consisting of the columns of $A$ indexed by $S$, $\beta_{S}$ denotes the corresponding subvector of $\beta$, and $\beta_{-S}$ denotes the subvector of $\beta$ indexed by the complement $S^{c}$. We adopt the standard conventions for zero-dimensional vectors and matrices. In particular, a $0\times n$ matrix and a $0\times 1$ vector are well defined, and the $0\times 0$ matrix is treated as a valid square matrix whose inverse is itself. Under these conventions, matrix multiplication involving zero-dimensional objects is well defined and dimensionally consistent. We work with the linear model
$y=X\beta_{0}+\varepsilon,$
where $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times p}$, and $\beta^{0}\in\mathbb{R}^{p}$. The noise vector $\varepsilon\in\mathbb{R}^{n}$ is arbitrary and no specific distributional assumption is imposed unless stated otherwise.
2.2 Lasso and Ridge: Definitions and Basic Properties
Definition 2.1 (Lasso estimator).
Let $\lambda_{\mathrm{L}}>0$ be a regularization parameter. The Lasso estimator $\hat{\beta}_{\mathrm{L}}\in\mathbb{R}^{p}$ is defined as
$\hat{\beta}_{\mathrm{L}}\in\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\lambda_{\mathrm{L}}\|\beta\|_{1}\right\}.$
Definition 2.2.
We define the key quantities associated with the Lasso:
  • The equicorrelation set:
    $E\coloneqq\{\,i\in\{1,2,\ldots,p\}:|\frac{1}{n}X_{i}^{\top}(y-X\hat{\beta}_{\mathrm{L}})|=\lambda_{\mathrm{L}}\,\}.$
  • Restricted Gram matrix:
    $\Sigma_{n,E}\coloneqq\frac{X_{E}^{\top}X_{E}}{n}.$
  • The equicorrelation signs:
    $s\coloneqq\operatorname{sign}({X_{E}}^{\top}(y-X\hat{\beta}_{\mathrm{L}})).$
Proposition 2.3.
The Karush–Kuhn–Tucker (KKT) conditions for the Lasso estimator can be written as follows; see, for example, (van de Geer, 2016):
$\frac{1}{n}X^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}\bar{z},$
where $\bar{z}\in\partial\|\hat{\beta}_{\mathrm{L}}\|_{1}$, whose $i$-th component is given by
$\bar{z}_{i}=\begin{cases}\mathrm{sign}((\hat{\beta}_{\mathrm{L}})_{i}),&\text{if }(\hat{\beta}_{\mathrm{L}})_{i}\neq 0,\\[6.0pt] u_{i}\in[-1,1],&\text{if }(\hat{\beta}_{\mathrm{L}})_{i}=0,\end{cases}\qquad\text{for }i=1,\dots,p.$
Moreover, it follows that
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Here, $\mathrm{sign}(\cdot)$ denotes the sign function. In addition, the KKT conditions on the active block imply
$\hat{\beta}_{\mathrm{L},E}=(X_{E}^{\top}X_{E})^{-1}\!\left(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s\right).$
Here, $(\hat{\beta}_{\mathrm{L}})_{i}$ denotes the $i$-th component of $\hat{\beta}_{\mathrm{L}}$, and for an index set $E$, $\hat{\beta}_{\mathrm{L},E}$ denotes the subvector restricted to $E$. The explicit expression for $\hat{\beta}_{\mathrm{L},E}$ holds when the Lasso solution is unique (see Proposition 2.6).
Assumption 2.4 (Column normalization).
The columns of the design matrix $X\in\mathbb{R}^{n\times p}$ are normalized as
$\|X_{j}\|_{2}^{2}=n,\qquad j=1,\dots,p.$
Assumption 2.5.
The columns of the design matrix $X$ are in general position. More precisely, for any integer $k<\min\{n,p\}$, no $k$-dimensional affine subspace of $\mathbb{R}^{n}$ contains more than $k+1$ elements of the signed set $\{\pm X_{1},\ldots,\pm X_{p}\}$, excluding antipodal pairs. Although the definition may appear somewhat technical, this condition is naturally satisfied when the entries of $X$ are drawn from a continuous distribution.
Proposition 2.6 (Uniqueness of the Lasso).
Under Assumption 2.5, the Lasso solution $\hat{\beta}_{\mathrm{L}}$ is unique, and its support is exactly $E$; that is, $\hat{\beta}_{j}\neq 0$ if and only if $j\in E$. Moreover, $\mathrm{null}(X_{E})=\{0\}$; see (Tibshirani, 2013).
Remark 2.7.
Proposition 2.6 ensures that the definition of the refined estimator is unambiguous. Throughout this paper, we impose Assumptions 2.4 and 2.5 without further mention.
Proposition 2.8 (Ridge regression estimator).
Let $\lambda^{\mathrm{r}}>0$ be a regularization parameter. The ridge regression estimator $\hat{\beta}^{\mathrm{r}}\in\mathbb{R}^{p}$ (Hoerl and Kennard, 1970) is defined by
$\hat{\beta}^{\mathrm{r}}=\arg\min_{\beta\in\mathbb{R}^{p}}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda^{\mathrm{r}}}{2}\|\beta\|_{2}^{2}\right\}.$
The Karush–Kuhn–Tucker (KKT) condition is given by
$\frac{1}{n}X^{\top}\bigl(y-X\hat{\beta}^{\mathrm{r}}\bigr)+\lambda^{\mathrm{r}}\hat{\beta}^{\mathrm{r}}=0,$
which implies the closed-form solution
$\hat{\beta}^{\mathrm{r}}=\bigl(X^{\top}X+n\lambda^{\mathrm{r}}I_{p}\bigr)^{-1}X^{\top}y.$
2.3 The Lasso–Ridge Estimator and Related Definitions
Definition 2.9.
A Lasso–Ridge estimator with $\lambda_{\mathrm{R}}\geq 0$ is defined as
$\hat{\beta}_{\mathrm{R}}=\underset{\beta\in\mathbb{R}^{p},\beta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-X\beta\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\beta-\hat{\beta}_{\mathrm{L}}\|_{2}^{2}\right\},$
or equivalently:
$\hat{\beta}_{\mathrm{R}}=\hat{\delta}+\hat{\beta}_{\mathrm{L}},$
$\quad\hat{\delta}=\underset{\delta\in\mathbb{R}^{p},\delta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-X\hat{\beta}_{\mathrm{L}}-X\delta\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta\|_{2}^{2}\right\}.$
Definition 2.10.
Let $\hat{\beta}_{\mathrm{L}}$ and $\hat{\beta}_{\mathrm{R}}$ denote the Lasso and Lasso–Ridge estimators of $\beta_{0}$, respectively. We define the prediction error gap:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}}):=\frac{1}{n}(\|X(\hat{\beta}_{\mathrm{L}}-\beta_{0})\|_{2}^{2}-\|X(\hat{\beta}_{\mathrm{R}}-\beta_{0})\|_{2}^{2}).$
Define the difference in prediction mean squared error:
$\mathrm{DMSE}:=\frac{1}{n}\,\mathbb{E}[\|X(\hat{\beta}_{\mathrm{L}}-\beta_{0})\|_{2}^{2}-\|X(\hat{\beta}_{\mathrm{R}}-\beta_{0})\|_{2}^{2}].$
A positive $\mathrm{DMSE}$ means $\hat{\beta}_{\mathrm{R}}$ strictly improves prediction MSE in expectation.
Definition 2.11 (Projection operators).
Let $A\in\mathbb{R}^{n\times k}$ be a matrix with full column rank, i.e., $\operatorname{rank}(A)=k$. The orthogonal projection onto the column space of $A$ is defined by
$P_{A}:=A(A^{\top}A)^{-1}A^{\top}.$
The projection onto the orthogonal complement of $\operatorname{col}(A)$ is
$P_{A}^{\perp}:=I_{n}-P_{A}.$
For any $u\in\mathbb{R}^{n}$, $P_{A}u$ is the projection of $u$ onto $\operatorname{col}(A)$, and $P_{A}^{\perp}u$ is the corresponding orthogonal residual. In particular, if $X_{E}\in\mathbb{R}^{n\times|E|}$ denotes the submatrix of $X$ with columns indexed by $E$, and $\operatorname{rank}(X_{E})=|E|$, then
$P_{X_{E}}$ $=X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top},$
$P_{X_{E}}^{\perp}$ $=I_{n}-P_{X_{E}}.$
3 Main Results
Theorem 3.1.
Consider our Lasso–Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have
$\mathrm{DMSE}$ $\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}\Bigl(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$\hskip 20.00003pt-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr)\Bigr\}.$
Here $|E|$ denotes the Lasso selected model size, which is random, so the tuning parameter $\lambda_{\mathrm{R}}$ is data-dependent.
Proof.
We start by rewriting the target term as follows:
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$ $=\frac{1}{n}\|X\beta_{0}-X\hat{\beta}_{\mathrm{L}}\|_{2}^{2}-\frac{1}{n}\|X\beta_{0}-X\hat{\beta}_{\mathrm{L}}-X\hat{\delta}\|_{2}^{2}$
$=\frac{2}{n}\langle X\beta_{0}-X\hat{\beta}_{\mathrm{L}},X\hat{\delta}\rangle-\frac{1}{n}\langle X\hat{\delta},X\hat{\delta}\rangle$
$=\frac{2}{n}\langle X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E},X_{E}\hat{\delta}_{E}\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$ (1)
Recalling the closed-form expression for the nonzero part of the Lasso estimator under the KKT conditions (Proposition 2.3), we have
$\hat{\beta}_{\mathrm{L},E}=(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s).$
The term $X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E}$ appearing in the previous step can be rewritten as follows:
$X\beta_{0}-X_{E}\hat{\beta}_{\mathrm{L},E}$ $=X\beta_{0}-X_{E}(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}y-n\lambda_{\mathrm{L}}s)$
$=X\beta_{0}-X_{E}(X_{E}^{\top}X_{E})^{-1}(X_{E}^{\top}(X\beta_{0}+\varepsilon)-n\lambda_{\mathrm{L}}s)$
$=P_{X_{E}}^{\perp}X\beta_{0}-P_{X_{E}}\varepsilon+n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,$
where $P_{X_{E}}$ denotes the orthogonal projection onto $\operatorname{col}(X_{E})$ and $P_{X_{E}}^{\perp}=I_{n}-P_{X_{E}}$ is the projection onto its orthogonal complement. Inserting this rewritten form into equality (1) and rearranging, we obtain
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$ $=\frac{2}{n}\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle-\frac{2}{n}\langle P_{X_{E}}\varepsilon,X_{E}\hat{\delta}_{E}\rangle$
$+\frac{2}{n}\langle n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,X_{E}\hat{\delta}_{E}\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$ (2)
Next, we deal with the four inner products appearing in the previous equation (2). First, note that for the first term, we have
$\langle P_{X_{E}}^{\perp}X\beta_{0},X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(I_{n}-X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
$=\hat{\delta}_{E}^{\top}(X_{E}^{\top}-X_{E}^{\top}X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})X\beta_{0}$
$=\hat{\delta}_{E}\mathbf{0}_{|E|\times n}X\beta_{0}$
$=0.$
Next, we simplify the second term in equation (2) using straightforward matrix multiplication:
$\langle P_{X_{E}}\varepsilon,X_{E}\hat{\delta}_{E}\rangle$ $=\hat{\delta}_{E}^{\top}X_{E}^{\top}(X_{E}(X_{E}^{\top}X_{E})^{-1}X_{E}^{\top})\varepsilon$
$=\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
We now simplify the third term:
$\frac{2}{n}\langle n\lambda_{\mathrm{L}}X_{E}(X_{E}^{\top}X_{E})^{-1}s,X_{E}\hat{\delta}_{E}\rangle$ $=2\lambda_{\mathrm{L}}\hat{\delta}_{E}^{\top}X_{E}^{\top}X_{E}(X_{E}^{\top}X_{E})^{-1}s$
$=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle.$
Substituting the above three equalities into equation (2), we obtain
$\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle-\frac{2}{n}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$ (3)
Consider the case where $E\neq\emptyset$. Then $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ admits the orthogonal decomposition:
$\Sigma_{n,E}=P\operatorname{diag}(\mu_{1},\dots,\mu_{|E|})P^{\top},$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}\geq 0$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the corresponding eigenvector matrix. According to Lemma A.2 in the Appendix, we obtain the following closed-form representation of $\hat{\delta}_{E}$:
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
By orthogonal decomposition, we have
$2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle$ $=2\lambda_{\mathrm{L}}^{2}s^{\top}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=2\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{1}{\mu_{1}+\lambda_{\mathrm{R}}},\dots,\frac{1}{\mu_{|E|}+\lambda_{\mathrm{R}}})P^{\top}s.$
Following a similar argument, we derive
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$ $=\frac{1}{n}\hat{\delta}_{E}^{\top}X_{E}^{\top}X_{E}\hat{\delta}_{E}$
$=\lambda_{\mathrm{L}}^{2}s^{\top}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}\Sigma_{n,E}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s.$
For notational convenience, define
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}}):=2\lambda_{\mathrm{L}}\langle\hat{\delta}_{E},s\rangle-\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle.$
Combining it with the previous two equations, we obtain the following simplified expression:
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}+2\lambda_{\mathrm{R}}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}+2\lambda_{\mathrm{R}}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s.$
Consider the function
$f_{1}(x)=\frac{x+2a}{(x+a)^{2}}.$
Notice that $\mu_{1}+\cdots+\mu_{|E|}=\operatorname{trace}(\Sigma_{n,E})=|E|$. Since $f_{1}(x)$ is monotone decreasing, for any $\mu_{1},\dots,\mu_{|E|}\leq|E|$,
$f_{1}(\mu_{i})\geq f_{1}(|E|),\quad i=1,\dots,|E|.$
Recall that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ where $c>2$. Since $s$ is a binary vector and $P$ is orthogonal (hence norm-preserving), the following inequalities hold:
$H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}+2c|E|}{(\mu_{1}+c|E|)^{2}},\dots,\frac{\mu_{|E|}+2c|E|}{(\mu_{|E|}+c|E|)^{2}})P^{\top}s$
$\geq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{|E|+2c|E|}{(|E|+c|E|)^{2}},\dots,\frac{|E|+2c|E|}{(|E|+c|E|)^{2}})P^{\top}s$
$=\frac{\lambda_{\mathrm{L}}^{2}}{|E|}\frac{2c+1}{(c+1)^{2}}\|P^{\top}s\|_{2}^{2}$
$=\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}.$
Recall that the above inequality holds on the event $\{E\neq\emptyset\}$. Since $H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})=0$ on $\{E=\emptyset\}$, we have
$\mathbb{E}[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})]$ $=\mathbb{E}\big[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E\neq\emptyset\}}+H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E=\emptyset\}}\big]$
$=\mathbb{E}[H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})\,\mathbf{1}_{\{E\neq\emptyset\}}]$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\,\mathbb{P}(E\neq\emptyset).$ (4)
Next, we consider the last remaining term in equation (3). From Lemma A.3 in the appendix, we know the following equation holds when $\lambda_{\mathrm{R}}>2|E|$:
$\|\hat{\delta}_{E}\|_{1}=s^{\top}\hat{\delta}_{E}.$
Since $\|\hat{\delta}_{E}\|_{1}=0$ on $\{E=\emptyset\}$, it is enough to consider the case $E\neq\emptyset$ for the upper bound:
$\|\hat{\delta}_{E}\|_{1}$ $=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}s^{\top}P\operatorname{diag}(\frac{\lambda_{\mathrm{R}}}{\mu_{1}+\lambda_{\mathrm{R}}},\dots,\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}})P^{\top}s$
$\leq\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}}\|P^{\top}s\|_{2}^{2}$
$\leq\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}\frac{\lambda_{\mathrm{R}}}{\mu_{|E|}+\lambda_{\mathrm{R}}}|E|$
$\leq\frac{\lambda_{\mathrm{L}}}{c}.$
Returning to the remaining term, the dual-norm inequality yields
$\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\mathbb{E}\langle\hat{\delta}_{E},X_{E}^{\top}\varepsilon\rangle$
$\leq\mathbb{E}[\|\hat{\delta}_{E}\|_{1}\,\|X_{E}^{\top}\varepsilon\|_{\infty}]$
$\leq\frac{\lambda_{\mathrm{L}}}{c}\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}].$ (5)
Next, we analyze the term $\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ in the following way:
$\|X_{E}^{\top}\varepsilon\|_{\infty}$ $=\|X_{E}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\subseteq T_{0}\}}+\|X_{E}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}$
$\leq\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\subseteq T_{0}\}}+\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}$
$\leq\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}+\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}.$
By taking expectations and applying the Cauchy–Schwarz inequality, we obtain
$\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ $\leq\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}\mathbf{1}_{\{E\not\subseteq T_{0}\}}]$
$\leq\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\not\subseteq T_{0}\}}^{2}]}$
$=\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}.$
Since the inequality holds for every subset $T_{0}\subseteq\{1,\dots,p\}$, we may take the infimum over all such $T_{0}$ and derive
$\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$ $\leq\inf_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]+\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr\}.$ (6)
Combining $(3)$, $(4)$, $(5)$, and $(6)$, we obtain the following inequalities, which conclude the proof:
$\mathrm{DMSE}$ $=\mathbb{E}\Delta(\hat{\beta}_{\mathrm{L}},\hat{\beta}_{\mathrm{R}})$
$=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\,\mathbb{P}(E\neq\emptyset)-\frac{2\lambda_{\mathrm{L}}}{cn}\mathbb{E}[\|X_{E}^{\top}\varepsilon\|_{\infty}]$
$\geq\sup_{T_{0}\subseteq\{1,\dots,p\}}\Bigl\{\frac{2\lambda_{\mathrm{L}}}{c}\Bigl(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}]$
$\hskip 20.00003pt-\frac{1}{n}\sqrt{\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}^{2}]}\sqrt{\mathbb{P}(E\not\subseteq T_{0})}\Bigr)\Bigr\}.$
Remark 3.2.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ under the column normalization of $X$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$ See Lemma A.4 in the appendix for a proof. Moreover, $\mathbb{P}(E\neq\emptyset)$, that is, the probability that the Lasso estimator is nonzero, is typically close to one under standard signal and tuning regimes. See Lemma A.5 and Remark A.6 in the appendix for a more detailed discussion.
Remark 3.3 (On the data-dependent choice of $\lambda_{\mathrm{R}}$).
This tuning choice $\lambda_{\mathrm{R}}=c|E|$ is adopted primarily to derive a clean and easily interpretable lower bound for the prediction MSE difference. The constant $c>2$ is mainly technical and can be relaxed under stronger design conditions, for example when the design matrix has bounded column inner products (i.e., weaker mutual coherence or more nearly orthogonal columns). Other choices of $\lambda_{\mathrm{R}}$ are also possible and would lead to alternative formulations, typically involving probability inequalities or less transparent bounds.
Remark 3.4 (Approximate support containment).
Note that $\sqrt{\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]}$ has the same Gaussian-max scale as $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}]$, while the third term in Theorem 3.1 is further attenuated by the factor $\mathbb{P}(E\not\subseteq T_{0})$. This motivates assuming the existence of a deterministic index set $T_{0}$ such that
$\mathbb{P}(E\subseteq T_{0})\geq 1-\eta$
for some fixed $0<\eta<1$, which controls the negative contribution in the lower bound. Let $S_{0}=\{j:\beta_{0,j}\neq 0\}$ denote the true support. When $T_{0}=S_{0}$, the event $E\subseteq S_{0}$ corresponds to the Lasso having no false inclusions, which holds with high probability under mild conditions; see, for example, (Hastie et al., 2015, Theorem 11.3, p. 302).
The result cited above also illustrates that such containment statements are generally weaker than exact variable selection consistency. In particular, support containment may hold without requiring strong signal conditions on $\beta_{0}$, and the irrepresentable conditions can be relaxed when it is considered with respect to a deterministic superset rather than the exact true support. Under such formulations, the probability of containment failure may still decay exponentially with the sample size.
It is worth noting that, instead of relying solely on the Cauchy–Schwarz inequality, one may invoke Hölder’s inequality, which provides additional flexibility in controlling and potentially attenuating the negative contribution. Next, considering the case $T_{0}=\{1,\dots,p\}$, we obtain the following corollaries:
Corollary 3.5.
Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, where $c>2$, then we have
$\mathrm{DMSE}$ $\geq\frac{2\lambda_{\mathrm{L}}}{c}(\frac{c(2c+1)}{2(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\frac{1}{n}\mathbb{E}[\|X^{\!\top}\varepsilon\|_{\infty}]).$
Remark 3.6.
Consider the function
$f_{2}(x)=\frac{x(2x+1)}{(x+1)^{2}},\quad x\geq 0.$
The function $f_{2}(x)$ is monotone increasing, and it converges to $2$ as $x\to\infty$. Suppose the penalty level $\lambda_{\mathrm{L}}$ of the Lasso estimator satisfies
$\lambda_{\mathrm{L}}>\frac{\mathbb{E}[\|X^{\!\top}\varepsilon\|_{\infty}]}{n\mathbb{P}(E\neq\emptyset)}.$
Then there exists a constant $c$ for the tuning parameter $\lambda_{\mathrm{R}}=c|E|$ such that
$\mathrm{DMSE}>0.$
In particular, since $\mathbb{P}(E\neq\emptyset)$ is typically close to one under the considered regimes, the Lasso can be prediction-suboptimal within this refinement class once $\lambda_{\mathrm{L}}$ moderately exceeds the noise-max scale $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]/n$.
Remark 3.7 (Noise robustness).
The main results do not rely on independence or Gaussianity of the noise. In particular, martingale difference errors can be accommodated under mild conditions. See the extension part for details. For further discussions, we first recall the standard bound for $\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]$. We then discuss how the structure of the design matrix $X$ critically influences the behavior of this Gaussian maximum.
Lemma 3.8 (Expectation of Gaussian maxima).
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $X=[X_{1},\ldots,X_{p}]\in\mathbb{R}^{n\times p}$ with column norms $\|X_{j}\|_{2}\leq\sqrt{n}$ for all $j$. Then
$\frac{1}{n}\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]\leq\sigma\sqrt{\frac{2\log(2p)}{n}}.$
For more details, see, for example, (van de Geer, 2016, Section 17.4, p. 243).
Remark 3.9 (Gaussian process interpretation).
Define $Z_{j}:=n^{-1}X_{j}^{\top}\varepsilon$, $j=1,\dots,p$. Then $\{Z_{j}\}_{j=1}^{p}$ is a centered Gaussian process indexed by $I=\{1,\dots,p\}$, with canonical metric
$d(i,j)=(\mathbb{E}|Z_{i}-Z_{j}|^{2})^{1/2}=\frac{\sigma}{n}\|X_{i}-X_{j}\|_{2},\qquad i,j\in I.$
Moreover,
$n^{-1}\|X^{\top}\varepsilon\|_{\infty}=\sup_{j\in I}|Z_{j}|.$
Thus, the stochastic term may be analyzed using Gaussian process tools, with the sub-Gaussian increment structure determined by the metric geometry of the column set. In particular, this representation makes general tools such as the chaining bound stated later naturally applicable.
Lemma 3.10 (Two-step bound under clustering).
Let $X=[X_{1},\dots,X_{p}]\in\mathbb{R}^{n\times p}$. Assume that the index set $\{1,\dots,p\}$ is partitioned into $K$ clusters $\{\mathcal{C}_{\ell}\}_{\ell=1}^{K}$, and for each cluster fix a representative vector $v_{\ell}\in\mathbb{R}^{n}$ satisfying $\|v_{\ell}\|_{2}\leq\sqrt{n}$. Write $v_{r(j)}$ for the representative of the cluster containing $j$, and suppose the clusters are $\delta$-tight in the sense that
$\|X_{j}-v_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j,$
for some $\delta\in(0,1]$. Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Then
$\frac{1}{n}\mathbb{E}[\|X^{\top}\varepsilon\|_{\infty}]\leq\frac{\sigma}{\sqrt{n}}(\sqrt{2\log(2K)}+\delta\sqrt{2\log(2p)}).$
A proof is provided in Appendix B.
Remark 3.11 (Subspace covering and induced clustering).
Suppose that the normalized columns $\{X_{j}/\sqrt{n}\}_{j=1}^{p}$ lie in an $r$-dimensional subspace of $\mathbb{R}^{n}$. Under our column normalization assumption, these points lie on the unit sphere $S^{r-1}$ within that subspace. By a standard covering argument on this sphere (see, for example, (Vershynin, 2018, Corollary 4.2.13)), for any $\delta\in(0,1]$, there exist vectors $\{v_{\ell}\}_{\ell=1}^{K}$ in the subspace, each satisfying $\|v_{\ell}\|_{2}=\sqrt{n}$, with $K\leq(1+2/\delta)^{r}$, such that
$\|X_{j}-v_{r(j)}\|_{2}\leq\delta\sqrt{n},\qquad\forall j.$
This provides a structural condition under which the assumption of Lemma 3.10 holds with $K\lesssim(1/\delta)^{r}$.
Notably, exact low-dimensional containment of the columns is generally incompatible with Assumption 2.5. However, the same clustering conclusion continues to hold under approximate low-dimensional structure: if each column $X_{j}$ admits a decomposition into a dominant component lying in an $r$-dimensional subspace and a small residual component orthogonal to that subspace, then the dominant part can be approximated by representatives as above, and the clustering bound still follows by the triangle inequality.
This phenomenon is closely related to settings with highly correlated covariates, particularly when the design exhibits low-dimensional structure and irrelevant columns lie close to the span of relevant ones, as discussed in (Dalalyan et al., 2017). Our analysis complements empirical evidence favoring smaller tuning parameters. Next, we cite a generic chaining bound, which provides a more general framework and is closely related to the structural viewpoint developed in our analysis.
Lemma 3.12 (Generic chaining bound (Talagrand, 2021)).
Let $(T,d)$ be a metric space. Suppose that $\{X_{t}:t\in T\}$ is a centered process whose increments are sub-Gaussian with respect to $d$. Then there exists an absolute constant $C>0$ such that
$\mathbb{E}\sup_{t\in T}X_{t}\leq C\,\gamma_{2}(T,d),$
where
$\gamma_{2}(T,d)=\inf_{\{T_{k}\}}\sup_{t\in T}\sum_{k\geq 0}2^{k/2}\,d(t,T_{k}),$
the infimum being taken over all admissible sequences $\{T_{k}\}$, that is, $|T_{0}|=1$ and $|T_{k}|\leq 2^{2^{k}}$ for all $k\geq 1$. See also, for example, (Vershynin, 2018) for further details on chaining methods, including Dudley’s entropy integral inequality, which provides a closely related perspective.
Remark 3.13 (Intrinsic stochastic complexity and tuning calibration).
The preceding lemmas indicate that the structure of the design matrix influences the range of Lasso tuning parameters for which prediction suboptimality occurs. In particular, in regimes where the containment error $\mathbb{P}(E\not\subseteq T_{0})$ is small, the lower bound on the DMSE increases as the term $\mathbb{E}[\|X_{T_{0}}^{\top}\varepsilon\|_{\infty}/n]$ decreases, where $T_{0}$ denotes a high-probability containment set for the Lasso-selected support. This suggests that the main situations leading to pronounced prediction suboptimality can be summarized as follows:
  • a low noise level;
  • a sparse true support, and the probability of false inclusions by the Lasso is small, or more generally, the Lasso-selected support is contained in a relatively small index set with high probability;
  • strong column correlations, or more precisely, a design with significant geometric redundancy, where many directions are governed by a smaller intrinsic set, or, more delicately, by refined chaining-based control.
Consequently, the lower bound provides a finite-sample explanation for why the universal Lasso tuning level $\lambda\asymp\sigma\sqrt{(\log p)/n}$ may be overly conservative in regimes where the effective stochastic complexity is governed by a smaller containment set $T_{0}$ and fewer representative directions. Finally, we consider the case where the Lasso recovers the true support with high probability under symmetric noise and establish a complementary theorem.
Theorem 3.14.
Let $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$, and let $S_{0}$ denote the true support. Consider our Lasso-Ridge method. Select the tuning parameter of the second step as $\lambda_{\mathrm{R}}=c|E|$, then we have
$\mathrm{DMSE}$ $\geq\frac{\lambda_{\mathrm{L}}}{\sqrt{c}}(\frac{\sqrt{c}(2c+1)}{(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\sigma(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}).$
Remark 3.15.
This theorem shows that when the Lasso selects the true support with high probability, its prediction performance can be improved by a suitable refinement over a wide range of tuning parameters. Moreover, it is known that variable selection consistency may occur when the Lasso tuning parameter does not vanish; see, for example, (Lahiri, 2021). Besides, under standard sign-consistency conditions such as the irrepresentable condition together with a beta-min assumption, the model selection error probability $\mathbb{P}(E\neq S_{0})$ becomes exponentially small; see (Zhao and Yu, 2006). In such settings, the present theorem implies that prediction improvement can be achieved. For completeness, the proof is given in Appendix B.
Remark 3.16 (Support stability beyond exact recovery).
The theorem does not require exact support recovery, i.e., $\mathbb{P}(E=S_{0})\to 1$. It suffices that the Lasso-selected support coincides with a fixed index set $S^{\star}$ with high probability, for instance, the support obtained in the noiseless case. Consequently, the refinement step can exploit this support stability to yield prediction improvement.
4 Discussion and Possible Extensions
Extensions to $\ell_{1}$-based methods.
The analysis developed in this paper is not specific to the classical Lasso, but depends more broadly on structural features shared by $\ell_{1}$-based procedures. At a high level, the extension follows the same general template: identify an intrinsic active or equicorrelation-type set induced by the regularization, perform the refitting step on this set, and derive a prediction dominance inequality from the definition of the second step, after which the main DMSE framework can be applied on the resulting intrinsic set.
This perspective is expected to extend, at least formally, to other $\ell_{1}$-based methods such as the Dantzig selector (Candès and Tao, 2007) and the Adaptive Lasso (Zou, 2006). For the Dantzig selector, one may analogously introduce an intrinsic active set associated with the constraint structure, while for the Adaptive Lasso with fixed weights, the same geometric decomposition can be adapted to the weighted $\ell_{1}$ setting, leading to a corresponding equicorrelation-type structure. A complete extension to these methods, especially for a precise DMSE analysis, requires additional technical developments and is therefore left for future work. Next, we discuss extensions to martingale noise structures.
Martingale noise.
The results of this paper do not fundamentally rely on independence or Gaussianity of the noise. In particular, the framework also accommodates martingale difference noise under suitable moment or conditional variance assumptions. Consider the fixed-design linear regression model $y_{i}=x_{i}^{\top}\beta_{0}+\varepsilon_{i}$, $i=1,\dots,n$, where $X\in\mathbb{R}^{n\times p}$ is deterministic with normalized columns $\|X_{j}\|_{2}^{2}=n$ for all $j=1,\dots,p$, and $(\varepsilon_{i},\mathcal{F}_{i})_{i=1}^{n}$ is a martingale difference sequence satisfying $\mathbb{E}(\varepsilon_{i}\mid\mathcal{F}_{i-1})=0$ and $\mathbb{E}(\varepsilon_{i}^{2}\mid\mathcal{F}_{i-1})\leq\sigma^{2}$ almost surely. Assume further that the design obeys a mild regularity condition such as $\max_{i,j}|x_{ij}|\leq K$. Then, for any index set $T$ and each $j\in T$,
$(X_{T}^{\top}\varepsilon)_{j}=\sum_{i=1}^{n}x_{ij}\varepsilon_{i}$
is a weighted sum of martingale differences. Hence, by Theorem A1.1.6 in (Nishiyama, 2021, p. 219), $\mathbb{E}\|X_{T}^{\top}\varepsilon/n\|_{\infty}$ is of order $\sqrt{\log|T|/n}$, up to constants. Substituting such a maximal inequality into Theorem 3.1 yields a corresponding range of tuning parameters for which the Lasso may exhibit suboptimal prediction performance under martingale noise. Such settings arise naturally in heteroskedastic regression, econometrics, and time-series models with predictable conditional variances.
5 Conclusions
This paper revisits the tuning paradigm for the Lasso from the perspective of prediction risk dominance. In particular, we propose tuning ranges in which suboptimality arises; more specifically, the resulting predictor can be dominated in mean squared prediction error by refitting on the Lasso-selected support. From this perspective, the results provide a quantitative explanation for empirical observations that smaller tuning parameters may perform favorably in correlated designs and related settings, and that refitting on the selected support of the Lasso may improve prediction performance.
Although the Lasso serves as the primary reference point, the arguments rely mainly on the structure of $\ell_{1}$-regularization and the associated DMSE decomposition, rather than on features unique to a specific estimator. Extending the analysis to more general settings, such as nonlinear models, is left for future work. Developing new frameworks for the standard linear regression model is also an interesting direction.
Appendix
Appendix A Proofs of Technical Lemmas
Lemma A.1 (Neumann Serious Expansion).
Let $A\in\mathbb{R}^{d\times d}$ with $\|A\|_{\infty}<1$. Then we have
$(I-A)^{-1}=\sum_{k=0}^{\infty}A^{k}.$
For more details, see, for example, (Gentle, 2024, Section 3.11.6, p. 197).
Lemma A.2 (Closed-form Representation under the KKT Conditions).
When $E\neq\emptyset$, the nonzero entries of $\hat{\delta}$ admit a simple expression:
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{L}}}{\lambda_{\mathrm{R}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
Proof.
Following from the second definition of our refitting step, we have
$\hat{\delta}=\underset{\delta\in\mathbb{R}^{p},\delta_{-E}=0}{\arg\min}\left\{\frac{1}{2n}\|y-\hat{\beta}_{\mathrm{L}}-X_{E}\delta_{E}\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta_{E}\|_{2}^{2}\right\},$
So when $E\neq\emptyset$, $\hat{\delta}_{E}$ could be regarded as a simple ridge regression estimator:
$\hat{\delta}_{E}=\underset{\delta\in\mathbb{R}^{|E|}}{\arg\min}\left\{\frac{1}{2n}\|y-X\hat{\beta}_{\mathrm{L}}-X_{E}\delta_{E}\|_{2}^{2}+\frac{\lambda_{\mathrm{R}}}{2}\|\delta_{E}\|_{2}^{2}\right\}.$
Recall the closed from representation under the KKT conditions of a ridge regression model in Proposition 2.8, we derive
$\hat{\delta}_{E}=\frac{1}{n}(\frac{1}{n}X_{E}^{\top}X_{E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}}).$
From the KKT conditions of the Lasso step, we obtain the following result:
$\frac{1}{n}X_{E}^{\top}(y-X\hat{\beta}_{\mathrm{L}})=\lambda_{\mathrm{L}}s.$
Combining the previous two equations. Also replacing $X_{E}^{\top}X_{E}/n$ with $\Sigma_{n,E}$ and factoring $\lambda_{\mathrm{R}}$ out of the inverse completes the proof. ∎
Lemma A.3 (Alternative Form of $\|\hat{\delta}_{E}\|_{1}$).
Consider the case where $E\neq\emptyset$ and $\lambda_{\mathrm{R}}>2|E|$. We have $\|\hat{\delta}_{E}\|_{1}=s^{\top}\hat{\delta}_{E}$.
Proof.
Notice that, it suffices to show that $\operatorname{sign}(s)=\operatorname{sign}(\hat{\delta}_{E})$. According to lemma A.2, we have
$\hat{\delta}_{E}=\frac{\lambda_{\mathrm{R}}}{\lambda_{\mathrm{L}}}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s.$
Thus, the problem can be reformulated as follows:
$\operatorname{sign}((\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s)=\operatorname{sign}(s).$
Notice that $\|\Sigma_{n,E}\|_{\infty}\leq|E|$, since $\|\Sigma_{n,E}\|_{\infty}=\max_{i\in E}\sum_{j\in E}|(\Sigma_{n,E})_{ij}|$ and each $|(\Sigma_{n,E})_{ij}|\leq 1$. When $\lambda_{\mathrm{R}}>2|E|$, we have
$\|\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}\|_{\infty}<\frac{1}{2}.$
Applying Neumann series expansion (Lemma A.1), we get
$(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}=\sum_{k=0}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}=I_{|E|}+\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}.$
Thus, we obtain
$(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}}+I_{|E|})^{-1}s=s+(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s,$
and our target is reduced to showing that
$\operatorname{sign}(s+(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s)=\operatorname{sign}(s).$
If each entry of the second term on the left-hand side has absolute value less than $1$, then the preceding equality follows. Hence, it suffices to prove that
$\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s\|_{\infty}<1.$
Noting that $s$ is a binary vector, we have
$\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})s\|_{\infty}$ $\leq\|(\sum_{k=1}^{\infty}(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k})\|_{\infty}\leq\sum_{k=1}^{\infty}\|(-1)^{k}(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})^{k}\|_{\infty}$
$\leq\sum_{k=1}^{\infty}\|(\frac{\Sigma_{n,E}}{\lambda_{\mathrm{R}}})\|_{\infty}^{k}<\sum_{k=1}^{\infty}(\frac{1}{2})^{k}=1.$
Lemma A.4.
Suppose that $\varepsilon\sim N(0,\sigma^{2}I_{n})$ and $\|X_{j}\|_{2}^{2}=n$ for all $j$. Then $\mathbb{E}[\|X^{\top}\varepsilon/n\|_{\infty}^{2}]\leq\sigma^{2}(2\log(4p)+1)/n.$
Proof.
For each $j$,
$\mathbb{P}(|X_{j}^{\top}\varepsilon|\geq t)$ $\leq 2\exp(-t^{2}/(2\sigma^{2}n)).$
Define $M=\|X^{\top}\varepsilon\|_{\infty}$. Then
$\mathbb{P}(M\geq t)$ $\leq 2p\exp(-t^{2}/(2\sigma^{2}n)).$
Using $\mathbb{E}[M^{2}]=\int_{0}^{\infty}2t\,\mathbb{P}(M\geq t)\,dt$, let $t_{0}=\sigma\sqrt{n}\sqrt{2\log(4p)}$. Then
$\int_{0}^{t_{0}}2t\,\mathbb{P}(M\geq t)\,dt$ $\leq t_{0}^{2}=2\sigma^{2}n\log(4p),$
$\int_{t_{0}}^{\infty}2t\,\mathbb{P}(M\geq t)\,dt$ $\leq 4p\int_{t_{0}}^{\infty}t\exp(-t^{2}/(2\sigma^{2}n))\,dt$
$=4p\,\sigma^{2}n\int_{\log(4p)}^{\infty}e^{-u}\,du$
$=\sigma^{2}n.$
Combining these two inequalities, we have
$\mathbb{E}[M^{2}]$ $\leq\sigma^{2}n\,(2\log(4p)+1),$
which concludes the proof. ∎
Lemma A.5.
Assume that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Then
$\mathbb{P}(E=\emptyset)\leq\Phi\!\left(\frac{\lambda_{\mathrm{L}}-\|\Sigma_{n}\beta_{0}\|_{\infty}}{\sigma/\sqrt{n}}\right),$
where $\Sigma_{n}:=X^{\top}X/n$.
Proof.
By the KKT conditions, $\hat{\beta}_{\mathrm{L}}=0$ if and only if
$\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}.$
Hence
$\mathbb{P}(E=\emptyset)$ $=\mathbb{P}(\hat{\beta}_{\mathrm{L}}=0)$
$=\mathbb{P}\!\left(\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}\right).$
Let $j^{\star}$ satisfy
$|(\Sigma_{n}\beta_{0})_{j^{\star}}|=\|\Sigma_{n}\beta_{0}\|_{\infty}.$
Then
$\left\{\frac{1}{n}\,\|X^{\top}y\|_{\infty}\leq\lambda_{\mathrm{L}}\right\}\subseteq\left\{\frac{1}{n}\,|X_{j^{\star}}^{\top}y|\leq\lambda_{\mathrm{L}}\right\}.$
Write
$m:=(\Sigma_{n}\beta_{0})_{j^{\star}}=\frac{1}{n}X_{j^{\star}}^{\top}X\beta_{0},\qquad|m|=\|\Sigma_{n}\beta_{0}\|_{\infty}.$
Using $y=X\beta_{0}+\varepsilon$, we obtain
$\mathbb{P}(E=\emptyset)\leq\mathbb{P}\!\left(\left|\frac{1}{n}X_{j^{\star}}^{\top}\varepsilon+m\right|\leq\lambda_{\mathrm{L}}\right).$
Let
$Z:=\frac{1}{n}X_{j^{\star}}^{\top}\varepsilon.$
Since $Z\sim N(0,\sigma^{2}/n)$ is symmetric about zero, we have $Z\stackrel{{\scriptstyle d}}{{=}}-Z$. Consequently,
$\mathbb{P}(|Z+m|\leq\lambda_{\mathrm{L}})$ $=\mathbb{P}(|-Z+m|\leq\lambda_{\mathrm{L}})$
$=\mathbb{P}(|Z-m|\leq\lambda_{\mathrm{L}}).$
If $m\geq 0$, then |Zm|=|Z|m|||Z-m|=|Z-|m||; if $m<0$, then m=|m|-m=|m| and |Zm|=|Z+|m|||Z-m|=|Z+|m||. In either case,
$\mathbb{P}(|Z+m|\leq\lambda_{\mathrm{L}})=\mathbb{P}(|Z+|m||\leq\lambda_{\mathrm{L}}).$
Because $|m|\geq 0$, the event $\{|Z+|m||\leq\lambda_{\mathrm{L}}\}$ is contained in $\{Z\leq\lambda_{\mathrm{L}}-|m|\}$, and therefore
$\mathbb{P}(E=\emptyset)\leq\mathbb{P}\!\left(Z\leq\lambda_{\mathrm{L}}-|m|\right)=\Phi\!\left(\frac{\lambda_{\mathrm{L}}-\|\Sigma_{n}\beta_{0}\|_{\infty}}{\sigma/\sqrt{n}}\right).$
Remark A.6.
In particular, if $\|\Sigma_{n}\beta_{0}\|_{\infty}\geq\lambda_{\mathrm{L}}+t\,\sigma/\sqrt{n}$ for some $t>0$, then
$\mathbb{P}(E=\emptyset)\leq\Phi(-t),$
which decays exponentially in $t^{2}$. Hence, under the usual scaling in which $\lambda_{\mathrm{L}}\to 0$ while $\|\Sigma_{n}\beta_{0}\|_{\infty}$ remains of constant order, the margin $\|\Sigma_{n}\beta_{0}\|_{\infty}-\lambda_{\mathrm{L}}$ is bounded away from zero, and consequently the probability that the Lasso solution is exactly zero is exponentially small.
Appendix B Proofs of Auxiliary Results
Proof of Lemma 3.10
Proof.
Recall that $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and, for each $j\in\{1,\dots,p\}$,
$Z_{j}=\frac{1}{n}X_{j}^{\top}\varepsilon.$
For each $j$, decompose
$|Z_{j}|=|\frac{1}{n}v_{r(j)}^{\top}\varepsilon+(Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon)|\leq|\frac{1}{n}v_{r(j)}^{\top}\varepsilon|+|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|.$
Taking maxima and expectations,
$\mathbb{E}\max_{j}|Z_{j}|\leq\mathbb{E}\max_{\ell\leq K}|\frac{1}{n}v_{\ell}^{\top}\varepsilon|+\mathbb{E}\max_{j\leq p}|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|.$
First, each $v_{\ell}^{\top}\varepsilon/n$ is centered Gaussian with variance
$\frac{\sigma^{2}}{n^{2}}\|v_{\ell}\|_{2}^{2}\leq\frac{\sigma^{2}}{n}.$
Applying Lemma 3.8 with $p$ replaced by $K$, we obtain
$\mathbb{E}\max_{\ell\leq K}|\frac{1}{n}v_{\ell}^{\top}\varepsilon|\leq\frac{\sigma}{\sqrt{n}}\sqrt{2\log(2K)}.$
Second, for each $j$,
$Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon=\frac{1}{n}(X_{j}-v_{r(j)})^{\top}\varepsilon,\quad\operatorname{Var}(Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon)\leq\frac{\sigma^{2}}{n}\delta^{2}.$
Thus $\{Z_{j}-v_{r(j)}^{\top}\varepsilon/n\}_{j=1}^{p}$ are centered Gaussian variables with variance at most $\sigma^{2}\delta^{2}/n$. By Lemma 3.8,
$\mathbb{E}\max_{j\leq p}|Z_{j}-\frac{1}{n}v_{r(j)}^{\top}\varepsilon|\leq\frac{\sigma}{\sqrt{n}}\,\delta\,\sqrt{2\log(2p)}.$
Combining the two bounds completes the proof. ∎
Proof of Theorem 3.14
Proof.
We first recall the following expression for the DMSE:
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\mathbb{P}(E\neq\emptyset)-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$ (1)
Also, recall that according to Lemma A.2, we have the closed-form expression for the Lasso–Ridge estimator:
$\hat{\delta}_{E}=\lambda_{\mathrm{L}}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s,$
which is determined once the values of $s$ and $|E|$ are fixed. We consider an alternative approach for the following term, with the tuning parameter $\lambda_{\mathrm{R}}$ in the second step set to $c|E|$:
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]+\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
$=\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E=S_{0}\}}]+\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}],$
where we adopt the same notation for any index set. In particular, $\Sigma_{n,S_{0}}\coloneqq X_{S_{0}}^{\top}X_{S_{0}}/n$, and $s_{0}\coloneqq\operatorname{sign}(\beta_{0,S_{0}})$ denotes the (deterministic) sign vector on the event $E=S_{0}$. Note that the vector $\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}$ is non-random; hence, by symmetry of the Gaussian noise $\varepsilon$, we have
$\frac{1}{n}\mathbb{E}\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle=0,$
Substituting this into the previous equation, we obtain
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]-\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}].$
First, we prove an upper bound of $\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$. Since $\|X_{E}\hat{\delta}_{E}\|_{2}^{2}=0$ on $\{E=\emptyset\}$, it is enough to consider the case $E\neq\emptyset$, where we have the following orthogonal decomposition based on the closed-form expression of the Lasso–Ridge estimator in Lemma A.2:
$\frac{1}{n}\|X_{E}\hat{\delta}_{E}\|_{2}^{2}$ $=\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$
$=\lambda_{\mathrm{L}}^{2}s^{\top}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}\Sigma_{n,E}(\Sigma_{n,E}+\lambda_{\mathrm{R}}I_{|E|})^{-1}s$
$=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+\lambda_{\mathrm{R}})^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+\lambda_{\mathrm{R}})^{2}})P^{\top}s,$
where $\mu_{1}\geq\mu_{2}\geq\cdots\geq\mu_{|E|}\geq 0$ denote the ordered eigenvalues of $\Sigma_{n,E}=X_{E}^{\top}X_{E}/n$ and $P$ is the corresponding eigenvector matrix. Consider the function
$f_{3}(x)=\frac{x}{(x+a)^{2}},\quad a>0.$
Note that $f_{3}(x)$ attains its maximum at $x=a$. Thus, we have
$\frac{\mu_{i}}{(\mu_{i}+c|E|)^{2}}\leq\frac{c|E|}{(c|E|+c|E|)^{2}},\quad i=1,\dots,|E|.$
Based on this, and recalling that we choose the tuning parameter $\lambda_{\mathrm{R}}=c|E|$, we derive
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle$ $=\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{\mu_{1}}{(\mu_{1}+c|E|)^{2}},\dots,\frac{\mu_{|E|}}{(\mu_{|E|}+c|E|)^{2}})P^{\top}s$
$\leq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{c|E|}{(c|E|+c|E|)^{2}},\dots,\frac{c|E|}{(c|E|+c|E|)^{2}})P^{\top}s$
$\leq\lambda_{\mathrm{L}}^{2}s^{\top}P\operatorname{diag}(\frac{c|E|}{(2c|E|)^{2}},\dots,\frac{c|E|}{(2c|E|)^{2}})P^{\top}s$
$\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c|E|}\|P^{\top}s\|_{2}^{2}$
$\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c}.$ (3)
Also note that, since $\|\varepsilon\|_{2}^{2}/\sigma^{2}\sim\chi^{2}_{n}$, we have
$\mathbb{E}\!\left[\|\varepsilon\|_{2}^{2}\right]=n\sigma^{2}.$
Now returning to our target, we have
$\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$ $\leq\frac{1}{n}\sqrt{\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\neq S_{0}\}}^{2}]}$
$=\frac{1}{n}\sqrt{\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{1}{n}\sqrt{\mathbb{E}[\|X_{E}\hat{\delta}_{E}\|_{2}^{2}\,\|\varepsilon\|_{2}^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\sqrt{\frac{\lambda_{\mathrm{L}}^{2}}{4c}\mathbb{E}[\frac{1}{n}\|\varepsilon\|_{2}^{2}]}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}\sqrt{\mathbb{P}(E\neq S_{0})}.$ (4)
Next, we derive an upper bound for
$\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
using a similar approach. First, note that on the event $E=S_{0}$, we have
$\frac{1}{n}\langle X_{E}\hat{\delta}_{E},X_{E}\hat{\delta}_{E}\rangle=\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}.$
Following an argument similar to that used to derive inequality (3), we obtain
$\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}\leq\frac{\lambda_{\mathrm{L}}^{2}}{4c}.$
Since $\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle$ is a Gaussian random variable, we have
$\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle^{2}]$ $=\sigma^{2}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}.$
Thus for the second term of our final target, we derive
$\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
$\leq\frac{1}{n}\sqrt{\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle^{2}]}\sqrt{\mathbb{E}[\mathbf{1}_{\{E\neq S_{0}\}}^{2}]}$
$\leq\frac{1}{n}\sqrt{\sigma^{2}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma}{\sqrt{n}}\sqrt{\frac{1}{n}\|\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0}\|_{2}^{2}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{cn}}\sqrt{\mathbb{P}(E\neq S_{0})}.$ (5)
Combining (4) and (5), we have
$\frac{1}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle$ $=\frac{1}{n}\mathbb{E}[\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]-\frac{1}{n}\mathbb{E}[\langle\lambda_{\mathrm{L}}X_{S_{0}}(\Sigma_{n,S_{0}}+c|S_{0}|I_{|S_{0}|})^{-1}s_{0},\varepsilon\rangle\mathbf{1}_{\{E\neq S_{0}\}}]$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}\sqrt{\mathbb{P}(E\neq S_{0})}+\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{cn}}\sqrt{\mathbb{P}(E\neq S_{0})}$
$\leq\frac{\sigma\lambda_{\mathrm{L}}}{2\sqrt{c}}(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}.$
Substituting this into equation (1), we obtain
$\mathrm{DMSE}$ $=\mathbb{E}H_{E,s}(\lambda_{\mathrm{L}},\lambda_{\mathrm{R}})-\frac{2}{n}\mathbb{E}\langle X_{E}\hat{\delta}_{E},\varepsilon\rangle.$
$\geq\lambda_{\mathrm{L}}^{2}\frac{2c+1}{(c+1)^{2}}\mathbb{P}(E\neq\emptyset)-\frac{\sigma\lambda_{\mathrm{L}}}{\sqrt{c}}(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}$
$=\frac{\lambda_{\mathrm{L}}}{\sqrt{c}}(\frac{\sqrt{c}(2c+1)}{(c+1)^{2}}\lambda_{\mathrm{L}}\mathbb{P}(E\neq\emptyset)-\sigma(1+\frac{1}{\sqrt{n}})\sqrt{\mathbb{P}(E\neq S_{0})}).$
Remark B.1.
Consider the function
$f_{4}(x)=\frac{\sqrt{x}(2x+1)}{(x+1)^{2}},\qquad x\geq 0.$
A direct calculation shows that $f_{4}(x)$ increases on $0,(3+\sqrt{17})/4]$, decreases thereafter, and attains its maximum value $f_{4,\max}\approx 0.79$ at $x\approx 1.78$.
Acknowledgement
The author gratefully acknowledges Professor Yoichi Nishiyama for his valuable guidance, insightful comments, and technical assistance.